# Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics

### v1

@@ 4.2 Moral Foundations Classification (MFC)
Table 2 presents the prompting questions of the inferential steps used in the MFC task. The input here is a simple RoT and the definitions of six MFs [^4] The definition of MFs are collected from the benchmark Ziems et al. (2022).. The inference steps are built as such:
  • (1)
    Step 1 identifies the action encoded by the RoT Forbes et al. (2020). This step prepares the target (action) for moral evaluation.
  • (2)
    Step 2 elicits the outcome of the action, namely, the behavioral consequence. One element of pragmatic inference is social norms, the variable that is probably most important in this task, given that RoTs were provided here without context. Norms for moral evaluations feature the physical and emotional harm to others that an action may incur Huebner et al. (2010); Kádár and Márquez-Reiter (2015); Turiel (1983). Therefore, Step 2 asks for the behavioral consequence as the linguistic manifestation of the social norms. This step setting also aligns with prior NLP studies showing that outcomes help establish the link between an action and its moral judgment Lindner et al. (2017); Lindner and Möllney (2019).
  • (3)
    Step 3 connects the action and its consequence to relevant MFs by textualizing their metapragmatic links.
Table 3 reports performance on development set instances with 1–3 moral foundation labels. Instances with more than three labels are relatively rare. We also report the average performance for each experimental setting. Across all training dataset sizes, model variants, and numbers of moral foundation labels, our method consistently outperforms the baseline methods. Comparing the results of base and base+, it is evident that incorporating MFs is beneficial, which aligns with findings from previous studies Ziems et al. (2022); Liu et al. (2025a).
Interestingly, for each model size, increasing the training dataset does not always lead to performance gains, and for each dataset size, using larger models also does not always result in improved performance for both the baseline and our method. We believe this is because a training dataset of the same size is insufficient to modify the knowledge that larger LLMs have already acquired during pre-training compared to smaller models. A notable feature of our proposed pragmatic inference is that, unlike the baseline methods whose performance drops significantly as the number of underlying MFs increases, our method exhibits much smaller performance degradation.
@@ 4.3 Moral Judgment
Input: There is a Prompt-Reply pair: {prompt} {reply}.
The moral foundations underlying this Prompt-Reply are {moral foundations}.
Inference: (1) Explain the definition of moral foundations {moral foundations}. (2) What conclusion
can be drawn from the Reply? (3) The moral judgment of the Reply is {Judgment}. Explain how the
conclusion of the Reply upholds or violates {moral foundations}, with reference to their definitions.
Table 4: Prompting Questions for the Task of Moral Judgment. All content within {} is provided by the benchmark. Answer examples for these prompting questions are available in Appendix A.4.
@@
Table 5: Accuracy of the Moral Judgment Task Across Different Training Dataset Sizes and Model Sizes. The optimal performance is highlighted with a bold font.
Input: There are six moral foundations and their DEFINITIONS are:{DEFINITIONS}.
There is a Prompt-Reply pair: {prompt} {reply}.
Inference: (1) Based on the Prompt, what conclusion can be drawn from the Reply? (2) Why the
conclusion of the Reply is associated to the moral foundations {moral foundations} by referring to
their definitions? (3) The moral judgment of the Reply is {judgment}. Explain how the conclusion of
the Reply obeys or violates the moral foundations of the {moral foundation}.
Table 6: Prompting Questions for the Joint Task of Moral Foundations Classification and Moral Judgment. All content within {} is provided by the benchmark. The answer example is shown in Appendix A.5.
Table 4 presents the prompting questions used in the second pragmatic inference framework that we have designed for moral judgment. Different from the first MFC task, where RoTs are not contextualized, the input here consists of the prompt–reply pair along with the underlying MFs. In other words, we provide the specific context (i.e., prompt+reply) and examine whether LLMs can effectively make moral judgments. The inference steps of the second task are:
  • (1)
    Similar to Step 1 in the first task (i.e., specifying actions), the first step specifies the moral foundations and prepares them for moral evaluation.
  • (2)
    Because the second task involves the specific context of a situation (prompt+reply), Step 2 here focuses on the context-specific interpretations of the reply, by concluding the reply. For example, the reply of “some people are built that way”alone may give rise to multiple interpretations, e.g., attributing one’s mistake to their nature, discriminating certain types of people, or merely stressing an inherent part of some people. However, when being related to the prompt of “fat people having hard or soft stomach”, the conclusion is that “the variation of stomachs is natural to humans’ physical construct”, hence preventing one’s physical appearance from being judged. This step supports the conclusion-deriving process of pragmatic inference from context Elder (2024), and, at the same time, reconstructs the literal meanings by connecting reply to prompt, which has been found to scaffold pragmatic interpretations Chen and Brown (2024); Chen and Lee (2021); Chen and Wang (2025).
  • (3)
    Step 3 connects the conclusion to the relevant moral foundations by textualizing their metapragmatic links, which is very similar to Step 3 in the first task.
@@
Table 5 presents the performance on the moral judgment task. Again, our pragmatic inference framework (ours) consistently outperforms both baseline methods (base, base+) across different training dataset sizes. Similar to the MFs classification task, increasing data size does not necessarily improve the model performance. Since moral judgment is a three-class classification task, where random guessing yields an accuracy of approximately 0.333, the baseline methods perform worse than random guessing in most settings of dataset and model sizes. In contrast, baseline methods perform much better than random guessing on the MFs classification task. The difference indicates that the pragmatic gap varies in size between the two tasks. This is understandable as the involvement of context (prompt+reply) in the second task brings about more variability in interpretations than the close alignment of RoT and normative interpretations, e.g., the RoT  “It is good to keep in touch with friends” can rather be easily attributed to the relational norms between friends. Nevertheless, the different pragmatic gap sizes give rise to a crucial question of how LLMs can infer both of them in one task.
4.4 Joint Moral Foundations Classification and Moral Judgment
@@
Building on the promising results of the first two tasks, the third task teaches LLMs to jointly infer MFs from a given situation [^5] Please note that in our first task, moral foundations classification, we classify MFs from the given RoT. and make corresponding moral judgments (denoted as MFC-Judgment). Table 6 illustrates the prompting questions for our proposed pragmatic inference for this task. The input has taken one part each from the above two tasks, namely, the situation (Prompt+Reply) and the definitions of MFs. This is closer to the real-life setting where a speaker only hears the interaction and has some (implicit) knowledge of MFs. The inferential steps are designed as follows:
  • (1)
    Step 1 is taken from the second task, given the need for context-specific interpretations of the reply.
  • (2)
    However, unlike the second task, this task only offers the specific MFs in training, but neither in test nor in input. In other words, models must learn from the metapragmatic links made between a conclusion of reply and a given MF in the course of training. Afterwards it can, in the test stage, predict the MFs and then associate them with the conclusion obtained in Step 1.
  • (3)
    Step 3 follows the final inference step of the second task.
@@
Llama-1B 5000 .637 .602
10000 .645 .612
23500 .639 .604
Llama-3B 5000 .643 .604
10000 .639 .616
23500 .676 .621
Table 7: Accuracy Performance of the Moral Judgment Task and the Joint Task Across Different Training Dataset Sizes and Model Architectures. The moral reasoning objective of these two tasks is the same.
Table 7 introduces the performance of the joint task compared to the moral judgment task wherein the ground-truth MFs are provided. Given the same moral reasoning objective, LLMs perform slightly worse on the joint task than on the standalone moral judgment task; however, their judgment performance still exceeds that of the baseline methods reported in Table 5. These results demonstrate that the current pragmatic inference framework enables LLMs to automatically infer underlying MFs and make moral judgments that are competitive with those based on ground-truth MFs, demonstrating the effectiveness of our proposed pragmatic inference framework.
@@ 5 Analysis
In the previous sections, we demonstrated the effectiveness of our proposed pragmatic inference methods. In this section, we provide additional analyses to further characterize them. Specifically, we: (1) conduct intervention experiments showing that our pragmatic inference methods guide LLMs to avoid heuristics and instead rely on MFs when making moral judgments; (2) demonstrate that our methods introduce lower perplexity; and (3) present a detailed analysis of individual MFs, howing that while performance on majority MFs is strong, dataset imbalance remains a key bottleneck to achieving robust moral reasoning on rare MFs.
5.1 Dependence on Moral Foundations
Prior studies show that LLMs may not leverage the input of MFs or other morality-related explanations that are provided to them to make some accurate moral decisions, often making their learning restricted to spurious correlation between situation and reasoning objective. Shapira et al. (2024); Liu et al. (2025b); Sanchez-Bayona and Agerri (2025). Just as  Liu et al. (2025b) demonstrate, LLMs correct immoral outputs to moral ones, while neglecting the underlying moral issues provided to them. Therefore, we conduct an intervention experiment to examine whether our framework has facilitated LLMs’ effective use of MFs.
Specifically, we replace the MFs, that have been automatically selected by the models in the third task (Question (2) in Table 6), with the ground-truth MFs. To ensure that other steps are not influenced by the replacement, we first let the LLMs complete their inference process, and then merely replace the output of MFs with the ground-truth ones, and finally prompt the LLMs to complete the remaining inferential steps and make a moral judgment.. By doing so, we restrict the change solely to the replacement of MFs. Any performance gains can thus illuminate the effective use of MFs by LLMs.
Figure 1: Accuracy of the MFC-Judgment Task (the Same Third Task) With Our Intervention Experiment. Across three dataset sizes and models, our intervention experiments introduce apparent performance gains.
Figure 1 exhibits that the replacement with ground-truth MFs consistently improves performance compared to the experiment without this intervention. In other words, the current pragmatic inference method has indeed facilitated LLMs to make effective use of MFs. Given MFs are commonly grounded in various moral situations, the finding suggests the potential for the current method to be generalized to other moral tasks.
@@ 5.2 Language Modeling
Although fine-tuning can introduce forgetting in LLMs Kirkpatrick et al. (2017); Liu et al. (2024) and degrade their general capabilities, Liu et al. (2025a) emphasize that fine-tuning for moral reasoning, when not accounting for the pragmatic nature of morality, tends to yield higher perplexity. To further demonstrate the benefits of our proposed pragmatic inference methods, we calculate the Perplexity of the optimal fine-tuned checkpoints of baseline approaches and our proposed approach. We use the WikiText dataset Merity et al. (2016) to compute perplexity. Figure 2 reports perplexity results across different training dataset sizes and tasks. Models fine-tuned with our pragmatic inference methods consistently exhibit substantially lower perplexity compared to baseline approaches.
@@ 5.3 Moral Foundation-wise Analysis
Since MFs play an essential role in our pragmatic inference framework (see the above Section 5.1), we wonder whether and how the different distribution of them may affect the performance of models in the MFC-Judgment task. Figure 3 illustrates MFs-wise performance, highlighting a correlation between the number of situations of each MF in the training dataset and the eventual accuracy of moral judgment. Notably, care accounts for about 40% of the training situations, leading to higher model performance in test situations involving this foundation. In contrast, performance is much lower for test situations involving the sanctity foundation. These empirical results suggest a space for our methods to further improve model performance, if given sufficient and balanced annotation data.
@@ 6 Discussion
Pragmatic Inference for Moral Reasoning. Different from existing approaches that train models to map moral situations with or without moral foundations to moral judgment  Hendrycks et al. (2020); Forbes et al. (2020); Jiang et al. (2021); Ziems et al. (2022); Jiang et al. (2025), this paper proposes several possible bridges that effectively link the two. The bridge is built by entextualizing the often unsaid variables (e.g., social norms, metapragmatic evaluations) in pragmatic inference, hence closing the gap between distributional semantics leveraged by LLMs and moral reasoning. Including the inferential steps is found to significantly increase the model performance. The findings are in line with previous studies that develop other pragmatic inferential methods for toxic language detection tasks  Sap et al. (2020); Zhou et al. (2023); Chen and Wang (2025). In fact, the concept of entextualizing pragmatic inferential process may be widely applicable to other non-logical reasoning tasks, such as the understanding of metaphors, humor, and sarcasm, which LLMs have thus far been found deficient in Barattieri di SanPietro et al. (2023); Ruis et al. (2023). Textualzing the inference processes for various pragmatic tasks also help to reveal the types of pragmatic knowledge that can be taught to LLMs and how they can be incorporated into model training.
Moral Foundations Theory for Moral Reasoning. We should note that it is not always uncontroversial to involve MFs in the pragmatic inference, because: (1) there is criticism that moral foundations lack neurological evidence for its role in moral reasoning and may oversimplify how humans perform moral reasoning Haidt and Joseph (2011); Braicovich (2022); and (2) they cannot capture cultural differences in moralities Graham et al. (2009, 2011b). While the latter still awaits for examinations with mutilingual datasets, our study, nevertheless, finds the effectiveness of MFs in machine learning, despite their drawbacks in human reasoning. This leaves a further question to future studies as to whether there is a MFs-free inferential pathway that allows LLMs to sufficiently emulate or even precede humans in moral reasoning.

### v2

@@ 4.2 Moral Foundations Classification (MFC)
Figure 1 presents the prompting questions of the inferential steps used in the MFC task. The input here is a simple RoT and the definitions of six MFs. The inference steps are built as such: (1) Step 1 identifies the action encoded by the RoT forbes2020social. This step prepares the target (action) for moral evaluation. (2) Step 2 elicits the outcome of the action, namely, the behavioral consequence. One element of pragmatic inference is social norms, the variable that is probably most important in this task, given that RoTs were provided here without context. Norms for moral evaluations feature the physical and emotional harm to others that an action may incur huebner_moral-conventional_2010; kadar_impoliteness_2015; Turiel1983-TURTDO. Therefore, Step 2 asks for the behavioral consequence as the linguistic manifestation of the social norms. This step setting also aligns with prior NLP studies showing that outcomes help establish the link between an action and its moral judgment lindner2017hera; lindner2019extracting. (3) Step 3 connects the action and its consequence to relevant MFs by textualizing their metapragmatic links.
Table 2 reports performance on development set instances with 1–3 moral foundation labels. Moral situations with more than three labels are relatively rare. We also report the average performance for each experimental setting. Across all training dataset sizes, model variants, and numbers of moral foundation labels, our method consistently outperforms the baseline methods. Comparing the results of base and base+, it is evident that incorporating MFs is beneficial, which aligns with findings from previous studies ziems2022moral; liu2025diagnosing.
Interestingly, for each model size, increasing the training dataset does not always yield performance gains. We believe one significant reason is that the additional data does not substantially expand the range of unique moral situations, and, as we discussed in the Section 1, moral reasoning depends on interpreting the implicatures of such situations taguchi2011teaching; taguchi2018contexts. On the other hand, for each dataset size, using larger models also does not always result in improved performance for both the baseline and our method. We believe this is because a training dataset of the same size is insufficient to modify the knowledge that larger LLMs have already acquired during pre-training compared to smaller models. A notable feature of our proposed pragmatic inference is that, unlike the baseline methods whose performance drops significantly as the number of underlying MFs increases, our method exhibits much smaller performance degradation.
@@ 4.3 Moral Judgment
In this section, we first introduce our inference method for making moral judgments based on ground-truth moral foundations. We then present an alternative inference method for jointly performing moral foundation classification and moral judgment. This joint approach guides LLMs to automatically identify the underlying moral foundations and subsequently make moral judgments based on the identified foundations.
4.3.1 Moral Judgment Based on Ground-truth Moral Foundations
Figure 2 presents the prompting questions used in the second pragmatic inference method that we have designed for moral judgment based on ground-truth MFs. Different from the first MFC task, where RoTs are not contextualized, the input here consists of the prompt–reply pair along with the underlying MFs. In other words, we provide the specific context (i.e., prompt+reply) and examine whether LLMs can effectively make moral judgments.
The inference steps of the second task are: (a) Step 1-3 is similar to the inference steps in the first task (i.e., specifying actions and consequences), but the difference is that we only care about actions that are relevant to the given MFs in the second task. (b) Step 4 connects the conclusion to the relevant moral foundations by textualizing their metapragmatic links. (c) Step 5 identifies the sentiment toward the actions mentioned in the Reply. We consider this step to be crucial for signaling the moral judgment of the Reply for two reasons. First, MFT provides a lens for establishing a meta-pragmatic link between the Reply and the moral judgment by evaluating how the actions uphold or violate specific MFs. Second, the expressed sentiment reflects the Reply’s stance toward actions that obey or violate these MFs. For example, if the Reply exhibits a positive stance toward an action that violates certain MFs, we infer a moral judgment of disagreement. (d) Step 6 aggregates the evidence from the preceding inference steps and aligns it with the ground-truth moral judgment label.
@@ 4.3.1 Moral Judgment Based on Ground-truth Moral Foundations
Table 3: Accuracy of the Moral Judgment Task with Ground-Truth Moral Foundations Across Training Dataset and Model Sizes. The optimal performance is highlighted with a bold font.
@@
Table 3 presents the performance on the moral judgment task with ground-truth MFs. Again, our pragmatic inference method (ours) consistently outperforms both baseline methods (base, base+) across different training dataset sizes. Similar to the MFC task, increasing data size does not necessarily improve the model performance. Since moral judgment is a three-class classification task, where random guessing yields an accuracy of approximately 0.333, the baseline methods perform worse than random guessing in some settings of dataset and model sizes. In contrast, baseline methods perform much better than random guessing on the MFC task. The difference indicates that the pragmatic gap varies in size between the two tasks. This is understandable as the involvement of context (prompt+reply) in the second task brings about more variability in interpretations than the close alignment of RoT and normative interpretations, e.g., the RoT  “It is good to keep in touch with friends” can rather be easily attributed to the relational norms between friends. Nevertheless, the different pragmatic gap sizes give rise to a crucial question of how LLMs can infer both of them in one task.
4.3.2 Joint Moral Foundations Classification and Moral Judgment
@@
Building on the promising results of the first two tasks, the third task teaches LLMs to jointly infer MFs from a given situation and make corresponding moral judgments (denoted as MFC-Judgment). Please note that in our first task, moral foundations classification, we classify MFs from the given RoT. For the prompting questions, we replace the step 3 and 4 in Figure 2 with the following prompt questions:
(3). Based on the consequences of these Actions, why are their underlying moral foundations classified as {moral_foundations} according to the provided DEFINITIONs?
(4). How do Actions influence (up-regulate or down-regulate) their underlying moral foundations based on the valence of their consequences—positive, negative, or neutral?
Therefore, we guide LLMs to learn to infer the underlying moral foundations, motivated by their strong performance on the MFC task (Figure 1). The input has taken one part each from the above two tasks, namely, the situation (Prompt+Reply) and the definitions of MFs. This is closer to the real-life setting where a speaker only hears the interaction and has some (implicit) knowledge of MFs.
@@
Llama-1B 5000 .708 .706
10000 .728 .718
23500 .710 .704
Llama-3B 5000 .712 .750
10000 .735 .748
23500 .710 .766
Mistral-7B 5000 .704 .694
10000 .760 .720
23500 .696 .704
Table 4: Accuracy Performance of the Moral Judgment Task and the Joint Task Across Different Training Dataset Sizes and Model Architectures. The moral reasoning objective of these two tasks is the same. The optimal performance is highlighted with a bold font.
MFC Judgment
Accuracy(#MFs=1) Accuracy(#MFs=2) Accuracy(#MFs=3) Average
Model Size CoT ours CoT ours CoT ours CoT ours CoT ours
Llama-1B 5000 .661 .890 .475 .856 .428 .806 .521 .851 .629 .706
10000 .677 .832 .545 .778 .468 .743 .564 .784 .652 .718
23500 .717 .770 .509 .796 .401 .743 .542 .770 .641 .704
Llama-3B 5000 .719 .851 .544 .822 .459 .798 .574 .824 .635 .750
10000 .700 .725 .507 .649 .441 .698 549 .691 .653 .748
23500 .727 .839 .533 .784 .414 .743 .558 .789 .665 .766
Mistral-7B 5000 .677 .803 .533 .634 .401 .577 .537 .671 .637 .694
10000 .653 .754 .515 .606 .523 .536 .563 .632 .628 .720
23500 .649 .768 .420 .604 .383 .559 .484 .643 .644 .704
Table 5: Performance on the Moral Foundations Classification and Moral Judgment using the CoT-driven versus our MFT-driven inference methods. The optimal performance is highlighted with a bold font. Please note, our reported performance on moral judgment (ours) is from the model trained through the joint task. Our MFT-driven approach outperforms CoT across all tasks, model sizes, and training data scales. The prompts used in these evaluations are shown in Appendix 7.
Table 4 reports the performance of the joint task in comparison to the moral judgment task where ground-truth MFs are provided. Given the same moral reasoning objective, i.e., making a moral judgment, LLMs achieve performance on the joint task that is competitive with the standalone moral judgment task. Notably, for the LLama-3B model, performance on the joint task surpasses that of the moral judgment task based on ground-truth MFs. Furthermore, performance on the joint task consistently exceeds that of the baseline methods reported in Table 3. These results demonstrate that the current pragmatic inference framework enables LLMs to automatically infer underlying MFs and make moral judgments that are competitive with those based on ground-truth MFs, demonstrating the effectiveness of our proposed pragmatic inference framework.
@@ 5 Analysis
In the previous sections, we demonstrated the effectiveness of our proposed pragmatic inference methods. In this section, we present additional analyses to further characterize these methods and highlight their benefits.
MFC Judgment
#MFs=1 #MFs=2 #MFs=3 Average
.694 .274 .147 .579 .466
Table 6: Performance of Direct LLM Prompting. The prompts used in these evaluations are provided in Appendix 6.
5.1 Performance of Direct LLM Prompting
Table 6 presents the results obtained by directly prompting the advanced reasoning model§§§deepseek-reasoning on the two tasks. Notably, its MFC performance is even lower than that of the Llama-3B baseline reported in Table 2, and its judgment performance falls below all baseline methods reported in Table 3. These results show that moral reasoning remains a challenging task even for advanced LLMs.
5.2 Effectiveness of Inference Driven by Moral Foundations Theory
Recall that in Section 4.1, we motivated the use of MFT to guide the inference process from two perspectives: linguistics and statistics. To empirically evaluate the effectiveness of MFT in our inference methods, we conduct additional experiments that generate training data by prompting off-the-shelf LLMs in a standard Chain-of-Thought (CoT) setting. We then compare the moral foundations classification and the moral judgment performance of models trained with our proposed MFT-driven inference methods against those trained with CoT-based inference.
Table 5 reports the experimental results comparing CoT-driven inference with our proposed MFT-driven approach. Across all model sizes, training data scales, and both tasks, the MFT-driven inference consistently outperforms the CoT baseline. These findings highlight the effectiveness of incorporating MFT to guide the inference process.
@@ 5.3 Language Modeling
Although fine-tuning can introduce forgetting in LLMs kirkpatrick2017overcoming; liu2024towards and degrade their general capabilities, liu2025diagnosing emphasize that fine-tuning for moral reasoning, when not accounting for the pragmatic nature of morality, tends to yield higher perplexity. This is also verified in other morality-relevant tasks such as social stereotype mitigation liu2025diagnosingperformancetradeoffmoral. To further demonstrate the benefits of our proposed pragmatic inference methods, we calculate the Perplexity of the optimal fine-tuned checkpoints of baseline approaches and our proposed approach. We use the WikiText dataset merity2016pointer to compute perplexity. Figure 3 reports perplexity results across different training dataset sizes and tasks. Models fine-tuned with our pragmatic inference methods consistently exhibit substantially lower perplexity compared to baseline approaches.
@@ 5.4 Moral Foundation-wise Analysis
Since MFs play an essential role in our pragmatic inference methods, we wonder whether and how the different distribution of them may affect the performance of models in the MFC-Judgment task. Figure 4 illustrates MFs-wise performance, highlighting a correlation between the number of situations of each MF in the training dataset and the eventual accuracy of moral judgment. Notably, care accounts for about 40% of the training situations, leading to higher model performance in test situations involving this foundation. In contrast, performance is much lower for test situations involving the sanctity foundation. These empirical results suggest a space for our methods to further improve model performance, if given sufficient and balanced annotation data.
@@ 6 Discussion
Pragmatic Inference for Moral Reasoning. Different from existing approaches that train models to map moral situations with or without moral foundations to moral judgment  hendrycks2020aligning; forbes2020social; jiang2021can; ziems2022moral; jiang2025investigating, this paper proposes several possible bridges that effectively link the two. The bridge is built by textualizing the often unsaid variables (e.g., social norms, metapragmatic evaluations) in pragmatic inference, hence closing the gap between distributional semantics leveraged by LLMs and moral reasoning. Including the inferential steps is found to significantly increase the model performance. The findings are in line with previous studies that develop other pragmatic inferential methods for toxic language detection tasks  sap2020social; zhou-etal-2023-cobra; chen2025pragmatic. In fact, the concept of entextualizing pragmatic inferential process may be widely applicable to other non-logical reasoning tasks, such as the understanding of metaphors, humor, and sarcasm, which LLMs have thus far been found deficient in barattieri_di_sanpietro_pragmatic_2023; ruis_goldilocks_2023. Textualzing the inference processes for various pragmatic tasks also help to reveal the types of pragmatic knowledge that can be taught to LLMs and how they can be incorporated into model training.
Moral Foundations Theory for Moral Reasoning. We should note that it is not always uncontroversial to involve MFs in the pragmatic inference, because: (1) there is criticism that moral foundations lack neurological evidence for its role in moral reasoning and may oversimplify how humans perform moral reasoning haidt2011moral; braicovich2022under; and (2) they cannot capture cultural differences in moralities graham2009liberals; graham2011mapping. While the latter still awaits for examinations with mutilingual datasets, our study, nevertheless, finds the effectiveness of MFs in machine learning reinig2024survey, despite their drawbacks in human reasoning. Additionally, there is also neurological evidence linking human neural responses to moral foundations in the process of moral judgment hopp2023moral. This leaves a further question to future studies as to whether there is a MFs-free inferential pathway that allows LLMs to sufficiently emulate or even precede humans in moral reasoning.